# Tight Bounds for Schrödinger Potential Estimation in Unpaired Data Translation

### v2

@@ 2 Main Result
We have to introduce the requirement ${\mathcal{T}}_{\infty}\varphi=0$ for all $\varphi\in\Phi\cup\{\varphi^{*}\}$ because of the fact that Schrödinger log-potentials are defined up to an additive constant. Thus, the condition ${\mathcal{T}}_{\infty}\varphi=0$ helps us to avoid ambiguity. Assumption 4, together with Assumptions 2 and 3, ensures that the random variables $\varphi(Y_{i})$, $1\leqslant i\leqslant n$, and $\log{\mathcal{T}}_{T}[e^{\varphi}](Z_{j})$, $1\leqslant j\leqslant n$, are sub-exponential. This allows us to quantify deviations of the empirical risk $\widehat{\mathcal{L}}(\varphi)$ (see (12)) from its expectation using concentration inequalities. In (Puchkin et al., 2025), the authors provided several examples of reference classes satisfying Assumption 4 including classes of ReLU neural networks and classes of logarithms of Gaussian mixtures. We will use the last ones in Section 3 with numerical experiments. According to (Korotin et al., 2024, Theorem 3.4), Gaussian mixtures possess a uniform approximation property. For this reason, a proper choice of a Gaussian mixtures class may lead to convincing practical results.
@@
We also suppose that the log-potential $\varphi^{*}$ associated with the optimal coupling $\pi^{*}$ meets Assumption 4 as well. While the requirement $\varphi^{*}(x)\leqslant M$ is mild, the lower bound $\varphi^{*}(x)\geqslant-L\|\Sigma^{-1/2}(x-m)\|^{2}-M$ needs a discussion. We are going to provide a sufficient condition for Assumption 4 to hold. Let us define $\varphi_{\max}^{*}=\sup\limits_{x\in\mathbb{R}^{d}}\varphi^{*}(x)<+\infty$. In view of (8), we have
@@
recently considered by Yakovlev and Puchkin (2025) in the context of generative diffusion models. Here $g:[0,1]^{p}\rightarrow\mathbb{R}^{d}$ is a smooth map defined on a low-dimensional latent space. In (Yakovlev and Puchkin, 2025), the authors discussed that such a model is quite general and suitable for description of high-dimensional data occupying a vicinity of a smooth low-dimensional manifold.
Finally, we assume that $\Phi$ is parametrized by a finite-dimensional parameter $\theta\in\mathbb{R}^{D}$:
$\Phi=\left\{\varphi_{\theta}:\theta\in\Theta\right\},$
where $\Theta$ is a subset of a $D$-dimensional cube $1,1]^{D}$ and each function $\varphi_{\theta}$ maps $\mathbb{R}^{d}$ onto $\mathbb{R}$. We suppose that the parametrization is sufficiently smooth in the following sense.
Assumption 5.
There exists $\Lambda\geqslant 0$ such that
$\left|\varphi_{\theta}(x)-\varphi_{\theta^{\prime}}(x)\right|\leqslant\Lambda\left(1+\|x\|^{2}\right)\|\theta-\theta^{\prime}\|_{\infty}$
for all $\theta,\theta^{\prime}\in\Theta$ and all $x\in\mathbb{R}^{d}$.
A similar condition appeared in (Puchkin et al., 2025), where the authors argued that Assumption 5 holds for the aforementioned classes of neural networks with ReLU activations and classes of logarithms of Gaussian mixtures. We refer a reader to (Puchkin et al., 2025, Section 4) for the details.
We are ready to formulate the main result of this section.
Theorem 1.
Then there exists
$T_{0}\lesssim\frac{1}{b}\log\log\frac{1}{\Upsilon(n,1)}\vee\frac{1}{b}\log\left(d+LR^{2}+\frac{Ld^{2}}{b}+M\right)$
$\quad+\Upsilon(n,\delta)\left(1\vee\log\frac{(1\vee L/b)(M+d)}{\Delta}\right),$
To our knowledge, Theorem 1 provides the first non-asymptotic high-probability upper bound on the KL-divergence between the optimal coupling $\pi^{*}$ and the empirical risk minimizer $\widehat{\pi}$. The previous result of Korotin et al. (2024) (see Theorem A.1 in their paper) concerned only the expected excess risk $\mathbb{E}{\mathsf{KL}}(\pi^{*},\widehat{\pi})-\Delta$. We also would like to note that our rate of convergence is sharper than the ${\mathcal{O}}(1/\sqrt{n})$ bound of Korotin et al. (2024) when the approximation error $\Delta$ is small. This is the case when there exists $\varphi\in\Phi$ approximating $\varphi^{*}$ with a good accuracy with respect to $L_{1}(\rho_{T})$- and $L_{1}({\mathcal{N}}(\mu,\Sigma))$-norms. Indeed, according to Lemma F.2 and the representation (9), we have
The main limitation of Theorem 1 is the requirement $T\gtrsim\log\log n$. It will become clear from Section 4 that it plays an essential role in the proof of Theorem 1 and is necessary for verification of a Bernstein-type condition. The assumption $T\gtrsim\log\log n$ arises from the structure of the Schrödinger problem: when $T$ is too small, the reference OU process does not have enough time to reach stable regime. Nevertheless, we would like to emphasize that the OU process is exponentially ergodic while the condition on $T$ exhibits the doubly logarithmic dependence. Hence, there is still a significant correlation between the initial and final points of the SDE governed stochastic process. In the context of image-to-image translation, this means that the style of initial images will be successfully transferred to the final ones.
3 Numerical Experiments
This section is devoted to numerical experiments illustrating the advantage of the Ornstein-Uhlenbeck dynamics (4) as the reference process. For this purpose, we adopt ideas behind the light Schrödinger bridge (LightSB) algorithm from (Korotin et al., 2024), where the authors suggested a simple yet quite efficient approach for generative modelling. Throughout this section, we use the notations $\sigma_{t}=(1-e^{-2bt})/(2b)$ and $m_{t}(x)=e^{-bt}x+(1-e^{-bt})m$, $0\leqslant t\leqslant T$. In what follows, we consider the process (4) with $T=1$ and $\Sigma=\varepsilon I_{d}$, where $\varepsilon>0$. The light Schrödinger bridge considers a reference class of log-potentials
$\Phi=\{\varphi_{\theta}:\theta\in\Theta\},$
such that the functions $v_{\theta}(y)=e^{\varphi_{\theta}(y)-\|y\|^{2}/(2\varepsilon\sigma_{1}^{2})}$, referred to as adjusted potentials, are nothing but Gaussian mixtures:
$v_{\theta}(y)=\sum_{k=1}^{K}\alpha_{k}\,\mathsf{p}(y;r_{k},\varepsilon S_{k}).$
Here, for any $k\in\{1,\dots,K\}$, $\mathsf{p}(\cdot\,;r_{k},\varepsilon S_{k})$ is the density of the Gaussian distribution with mean $r_{k}$ and covariance $\varepsilon S_{k}$. In this case, the parameter $\theta=\{(\alpha_{k},r_{k},S_{k}):1\leqslant k\leqslant K\}$ is just the set of triplets. The main advantage of such parametrization is that the integral
${\mathcal{T}}_{1}[e^{\varphi_{\theta}}](x)=\int\limits_{\mathbb{R}^{d}}e^{\varphi_{\theta}(y)}\,\mathsf{q}_{1}(y\,|\,x)\,{\mathrm{d}}y$
admits a closed-form expression. This significantly saves computational resources. At the same time, according to (Korotin et al., 2024, Theorem 3.4), choosing a proper class of Gaussian mixtures, one can approximate quite complex potentials. With the described $\varphi_{\theta}$, the corresponding coupling $\pi^{\varphi_{\theta}}$ and the risk ${\mathcal{L}}(\varphi_{\theta})$ the (see (7) and (10)) reduce to
$\pi^{\varphi_{\theta}}(x,y)=\frac{\rho_{0}(x)\,e^{y^{\top}m_{1}(x)/(\varepsilon\sigma_{1}^{2})}v_{\theta}(y)}{c_{\theta}(x)}$
$c_{\theta}(x)=\int\limits_{\mathbb{R}^{d}}e^{y^{\top}m_{1}(x)/(\varepsilon\sigma_{1}^{2})}v_{\theta}(y)\,{\mathrm{d}}y=\sum\limits_{k=1}^{K}\widetilde{\alpha}_{k}(x),\quad\widetilde{\alpha}_{k}(x)=\alpha_{k}\exp\left\{\frac{r_{k}^{\top}m_{1}(x)}{\varepsilon\sigma_{1}^{2}}+\frac{\|S_{k}^{1/2}m(x)\|^{2}}{2\varepsilon\sigma_{1}^{2}}\right\},$
and $r_{k}(x)=r_{k}+S_{k}x$. The log-potential estimate $\widehat{\varphi}$ is constructed through minimization of the empirical risk
over all admissible values of $\theta$. We call the described algorithm LightSB-OU and provide its pseudocode in Appendix G.2. The standard LightSB algorithm Korotin et al. (2024) performs the same steps but replaces $\mathsf{q}_{1}(y\,|\,x)$ with the Gaussian kernel $(2\pi\varepsilon)^{-d/2}e^{-\|y-x\|^{2}/(2\varepsilon)}$ corresponding to the transition density of the scaled Wiener process (3) with $\sigma=\sqrt{\varepsilon}$.
Below we thoroughly evaluate our solver on setups with both synthetic (Section 3.1) and real data distributions (Sections 3.2 and 3.3). The method demonstrates a stable increase in quality relative to the usual LightSB and shows itself well in comparison with other Schrödinger bridge solvers. All code and notebooks with experiments are publicly available at GitHub.
3.1 Evaluation on the Gaussian Mixture
We also evaluated our modified LightSB algorithm on a Gaussian mixture problem. For this experiment, we used mixtures of 25 Gaussians under three configurations: uniform grid with standard covariance matrices, random grid locations with standard covariance matrices, uniform grid with anisotropic random covariance matrices, with the sliced approximation of the Wasserstein distance $\mathbb{W}_{1}$ as the metric, see Appendix G.2 for definition of metrics.
Table 1 summarizes the performance for $K=30$ potential approximation components. With a sufficient number of potential approximation components, our approach demonstrates consistent improvements in both $\mathbb{W}_{1},$ MMD and Mode Coverage (see Appendix G.3 for detailed plots and Appendix G.5 for hyperparameters values used).
Experiment Metric Value
LightSB LightSB-OU
Standard $\text{Sliced }\mathbb{W}_{1}$ 0.26 $\pm$ 0.016 0.14 $\pm$ 0.021
MMD 0.0004 $\pm$ 0.0001 0.0003 $\pm$ 0.0001
Mode Coverage 24.2 $\pm$ 0.4 25
Irregular $\text{Sliced }\mathbb{W}_{1}$ 0.52 $\pm$ 0.024 0.21 $\pm$ 0.029
MMD 0.0024 $\pm$ 0.0003 0.0004 $\pm$ 0.0002
Mode Coverage 21.8 $\pm$ 0.40 24.6 $\pm$ 0.49
Anisotropic $\text{Sliced }\mathbb{W}_{1}$ 0.25 $\pm$ 0.017 0.18 $\pm$ 0.019
MMD 0.0007 $\pm$ 0.0002 0.0006 $\pm$ 0.0002
Mode Coverage 24.4 $\pm$ 0.49 25
Table 1: Comparison of LightSB and LightSB-OU performance for $K=30$ across experiments for 25 Gaussians
3.2 Evaluation on the Single Cell Data
Solver $\mathbb{W}_{1}$ metric
OT-CFM (Tong et al., 2023a) $0.790\pm 0.068$
$\mathrm{SF}]^{2}$M-Exact (Tong et al., 2023b) $0.793\pm 0.066$
LightSB-OU (ours) $0.815\pm 0.016$
LightSB (Korotin et al., 2024) $0.823\pm 0.017$
Reg. CNF (Finlay et al., 2020) $0.825$
T. Net (Tong et al., 2020) $0.848$
DSB (De Bortoli et al., 2021) $0.862\pm 0.023$
I-CFM (Tong et al., 2023a) $0.872\pm 0.087$
$\mathrm{SF}]^{2}$M-Geo (Tong et al., 2023b) $0.879\pm 0.148$
NLSB (Koshizuka and Sato, 2022) $0.970$
$\mathrm{SF}]^{2}$M-Sink (Tong et al., 2023b) $1.198\pm 0.342$
SB-CFM (Tong et al., 2023a) $1.221\pm 0.380$
DSBM (Shi et al., 2023) $1.775\pm 0.429$
Table 2: The quality of intermediate distribution restoration of single-cell data by different methods.
For a more comprehensive comparison with LightSB, we conducted experiments on biological data (Tong et al., 2020). Following (Korotin et al., 2024), we formulated the problem as transporting the cell distribution at time $t_{i-1}$ to time $t_{i+1}$ for $i\in\{1,2,3\}$. We use results for other methods from (Tong et al., 2023b), whose authors were the first to consider this setup in (Tong et al., 2020). We then predicted the cell distribution at the intermediate time $t_{i}$ and computed the Wasserstein distance $\mathbb{W}_{1}$ between the predicted distribution and the ground truth. The results were averaged across all three setups ($i=1,2,3$), see Table 2.
To ensure statistical robustness, we repeated the experiment five times. Using the default LightSB parameters from the original paper, the baseline achieved $\mathbb{W}_{1}=0.823$, while our OU-modified version achieved $\mathbb{W}_{1}=0.815$. The hyperparameter values are reported in Appendix G.5.
3.3 Evaluation on Unpaired Image-to-Image Translation
Figure 1: Translation results from the latent space of ALAE using LightSB (left) and LightSB - OU(right). Top: Adult to Child — skin color and face shape preserved. Middle: Male to Female — glasses and face shape preserved. Bottom: Male to Female — skin color and face shape preserved.
To further demonstrate the advantages of our method, we evaluate it on an unpaired image-to-image translation task. Following the experimental setup of (Korotin et al., 2024), we employ the ALAE autoencoder (Pidhorskyi et al., 2020) and focus on the Adult-to-Child translation task using the full $1024\times 1024$ FFHQ dataset (Karras et al., 2019).
Below, we present comparative translation results for selected samples in Figure 1. While the baseline LightSB successfully performs the translation, it sometimes fails to preserve key attributes of the source distribution such as skin tone, eye color, and facial structure leading to noticeable deviations. In contrast, our model achieves more accurate translations while better retaining the distinctive features of the original distribution. We emphasize that these examples are not intended to suggest universal superiority but rather to showcase specific strengths of our method in certain cases (the hyperparameter values are specified in Appendix G.5).
4 Proof Insights
Finally, we would like to discuss some intuition behind the proof of Theorem 1. Rigorous derivations are deferred to Appendix A. As we discussed in Section 2, Assumptions 2, 3, and 4 allow us to use concentration inequalities for sub-exponential random variables and quantify the deviation of the empirical risks $\widehat{\mathcal{L}}(\varphi)$, $\varphi\in\Phi$, from its expectations. However, to get rates of convergence possibly sharper than ${\mathcal{O}}(1/\sqrt{n})$, we must verify a Bernstein-type condition. That is, we have to show that both

### v3

@@ 3 Main Result
We have to introduce the requirement ${\mathcal{T}}_{\infty}\varphi=0$ for all $\varphi\in\Phi\cup\{\varphi^{*}\}$ because of the fact that Schrödinger log-potentials are defined up to an additive constant. Thus, the condition ${\mathcal{T}}_{\infty}\varphi=0$ helps us to avoid ambiguity. Assumption 4, together with Assumptions 2 and 3, ensures that the random variables $\varphi(Y_{i})$, $1\leqslant i\leqslant n$, and $\log{\mathcal{T}}_{T}[e^{\varphi}](Z_{j})$, $1\leqslant j\leqslant n$, are sub-exponential. This allows us to quantify deviations of the empirical risk (13) from its expectation using concentration inequalities. In Puchkin et al. (2025), the authors provided several examples of reference classes satisfying Assumption 4, including classes of ReLU neural networks and classes of logarithms of Gaussian mixtures. We will use the last ones in Section 4 with numerical experiments. According to Korotin et al. (2024, Theorem 3.4), Gaussian mixtures possess a uniform approximation property. For this reason, a proper choice of a Gaussian mixtures class may lead to convincing practical results. Further discussion of Assumption 4 is deferred to Appendix A.
Finally, we assume that $\Phi$ is parametrized by a finite-dimensional parameter $\theta\in\mathbb{R}^{D}$: $\Phi=\left\{\varphi_{\theta}:\theta\in\Theta\right\}$, where $\Theta$ is a subset of a $D$-dimensional cube $1,1]^{D}$ and each function $\varphi_{\theta}$ maps $\mathbb{R}^{d}$ onto $\mathbb{R}$. We suppose that the parametrization is sufficiently smooth in the following sense.
Assumption 5.
The class $\Phi$ is parametric: $\Phi=\left\{\varphi_{\theta}:\theta\in\Theta\right\}$, where $\Theta\subseteq[-1,1]^{D}$. Moreover, there exists $\Lambda\geqslant 0$ such that
$\left|\varphi_{\theta}(x)-\varphi_{\theta^{\prime}}(x)\right|\leqslant\Lambda\left(1+\|x\|^{2}\right)\|\theta-\theta^{\prime}\|_{\infty}\quad\text{for all $\theta,\theta^{\prime}\in\Theta$ and all $x\in\mathbb{R}^{d}$.}$
A similar condition appeared in Puchkin et al. (2025), where the authors argued that Assumption 5 holds for the aforementioned classes of neural networks with ReLU activations and classes of logarithms of Gaussian mixtures. We refer a reader to Puchkin et al. (2025, Section 4) for the details.
We are ready to formulate the main result of this section.
Theorem 1.
Then there exists
$T_{0}\lesssim\frac{1}{b}\log\log\frac{1}{\Upsilon(n,1)}\vee\frac{1}{b}\log\left(d+LR^{2}+\frac{Ld^{2}}{b}+M\right)\lesssim\frac{\log\log n}{b}$
${\mathsf{KL}}(\pi^{*},\widehat{\pi})-\Delta$ $\lesssim\sqrt{\Upsilon(n,\delta)\Delta\left(1\vee\log\frac{(b\vee L)(M+d)}{b\Delta}\right)}$
$\quad+\Upsilon(n,\delta)\left(1\vee\log\frac{(b\vee L)(M+d)}{b\Delta}\right),$ (16)
Remark 2.
A reader can note that the quantity $\Upsilon(n,\delta)$ introduced in Theorem 1 scales as
$\Upsilon(n,\delta)\lesssim\big(\log n+\log d+\log(1/\delta)\big)\frac{Dd\log n}{n}.$
If the class $\Phi$ is rich enough in a sense that $\Delta\lesssim\Upsilon(n,\delta)$, then the upper bound (1) simplifies to
${\mathsf{KL}}(\pi^{*},\widehat{\pi})\lesssim\Upsilon(n,\delta)\log n\lesssim\big(\log n+\log d+\log(1/\delta)\big)\frac{Dd\log^{2}n}{n}.$
To our knowledge, Theorem 1 provides the first non-asymptotic high-probability upper bound on the KL-divergence between the optimal coupling $\pi^{*}$ and the empirical risk minimizer $\widehat{\pi}$. The previous result of Korotin et al. (2024) (see Theorem A.1 in their paper) concerned only the expected excess risk $\mathbb{E}{\mathsf{KL}}(\pi^{*},\widehat{\pi})-\Delta$. We also would like to note that our rate of convergence is sharper than the ${\mathcal{O}}(1/\sqrt{n})$ bound of Korotin et al. (2024) when the approximation error $\Delta$ is small. This is the case when there exists $\varphi\in\Phi$ approximating $\varphi^{*}$ with a good accuracy with respect to $L_{1}(\rho_{T})$- and $L_{1}({\mathcal{N}}(\mu,\Sigma))$-norms. Indeed, according to Lemma E.2 and the representation (10), we have
Theorem 1 requires $bT$ to satisfy the bound $bT\gtrsim\log\log n$. It will become clear from Section B.1 that it plays an essential role in the proof of Theorem 1 and is necessary for verification of a Bernstein-type condition. The assumption $bT\gtrsim\log\log n$ seems to be necessary for the fast rates, because of the loss function (see (13)). In the limiting case $T=0$ the loss becomes linear in $\varphi$ and has no curvature. Nevertheless, we would like to emphasize the following. First, the inequality $bT\gtrsim\log\log n$ should be considered as a condition on $bT$, rather than on the time horizon $T$ alone. This means that we can take $T=1$ (as it is usually done in Schrodinger bridge problems) and $b\gtrsim\log\log n$. Second, a violation of the condition $bT\gtrsim\log\log n$ does not mean that the ERM $\widehat{\pi}$ becomes inconsistent. Using the standard approach based on Rademacher complexities, one can show that the KL-divergence between $\pi^{*}$ and $\widehat{\pi}$ will still tend to zero but at a slower rate ${\mathcal{O}}(1/\sqrt{n})$. Finally, the OU process is exponentially ergodic while the condition on $bT$ exhibits the doubly logarithmic dependence. Hence, there is still a significant correlation between the initial and final points of the SDE-governed stochastic process. In the context of image-to-image translation, this means that the style of initial images will be successfully transferred to the final ones.
In generative diffusion models one is often interested in accuracy of the optimal drift estimation. According to Dai Pra (1991, Theorem 3.2), the optimally controlled process $X_{t}^{*}$, $0\leqslant t\leqslant T$, obeys the SDE
$X_{0}^{*}\sim\rho_{0},\quad{\mathrm{d}}X_{t}^{*}=-b(X_{t}^{*}-m){\mathrm{d}}t+u^{*}(X_{t}^{*},t){\mathrm{d}}t+\Sigma^{1/2}{\mathrm{d}}W_{t},\quad 0<t<T,$ (17)
where $u^{*}(x,t)=\Sigma\nabla\log{\mathcal{T}}_{T-t}[e^{\varphi^{*}}](x)$. With the estimate $\widehat{\varphi}$ at hand (see (12) for the definition), we can consider the corresponding process
$\widehat{X}_{0}\sim\rho_{0},\quad{\mathrm{d}}\widehat{X}_{t}=-b(\widehat{X}_{t}-m){\mathrm{d}}t+\widehat{u}(\widehat{X}_{t},t){\mathrm{d}}t+\Sigma^{1/2}{\mathrm{d}}W_{t},\quad 0<t<T,$ (18)
with $\widehat{u}(x,t)=\Sigma\nabla\log{\mathcal{T}}_{T-t}[e^{\widehat{\varphi}}](x)$. This brings us to a natural question whether $\widehat{u}(x,t)$ is close $u^{*}(x,t)$. According to Girsanov’s theorem (see, for example, Domingo-Enrich et al. (2025, Theorem 2 and eq. (138))) it holds that
[…]
The expectation in the left-hand side is taken with respect to the trajectory of $X_{t}^{*}$, $0<t<T$, conditioned on data $Y_{1},\dots,Y_{n}$ and $Z_{1},\dots,Z_{n}$. Hence, Theorem 1 also establishes a high-probability upper bound on the control estimation.
4 Numerical Experiments
This section is devoted to numerical experiments illustrating the advantage of the Ornstein-Uhlenbeck dynamics (4) as the reference process. For this purpose, we adopt ideas behind the light Schrödinger bridge (LightSB) algorithm from Korotin et al. (2024), where the authors suggested a simple yet quite efficient approach for generative modelling. We do not pursue the goal of developing a state-of-the-art method, it goes beyond the scope of the present paper. However, we would like to illustrate how a proper choice of the base process improves data generation. Throughout this section, we use the notations $\sigma_{t}^{2}=(1-e^{-2bt})/(2b)$ and $m_{t}(x)=e^{-bt}x+(1-e^{-bt})m$, $0\leqslant t\leqslant T$. In what follows, we consider the process (4) with $T=1$ and $\Sigma=\varepsilon I_{d}$, where $\varepsilon>0$. The light Schrödinger bridge considers a reference class of log-potentials $\Phi=\{\varphi_{\theta}:\theta\in\Theta\}$, such that the functions $v_{\theta}(y)=e^{\varphi_{\theta}(y)-\|y\|^{2}/(2\varepsilon\sigma_{1}^{2})}$, referred to as adjusted potentials, are nothing but Gaussian mixtures:
$v_{\theta}(y)=\sum_{k=1}^{K}\alpha_{k}\,\mathsf{p}(y;r_{k},\varepsilon\sigma_{1}^{2}S_{k}).$ (19)
Here, for any $k\in\{1,\dots,K\}$, $\mathsf{p}(\cdot\,;r_{k},\varepsilon\sigma_{1}^{2}S_{k})$ is the density of the Gaussian distribution with mean $r_{k}$ and covariance $\varepsilon\sigma_{1}^{2}S_{k}$. In this case, the parameter $\theta=\{(\alpha_{k},r_{k},S_{k}):1\leqslant k\leqslant K\}$ is just the set of triplets. The main advantage of such parametrization is that the Ornstein-Uhlenbeck transform ${\mathcal{T}}_{1}[e^{\varphi_{\theta}}](x)$ admits a closed-form expression. This significantly saves computational resources. At the same time, according to Korotin et al. (2024, Theorem 3.4), choosing a proper class of Gaussian mixtures, one can approximate quite complex potentials. With the described $\varphi_{\theta}$, the corresponding coupling $\pi^{\varphi_{\theta}}$ and the risk ${\mathcal{L}}(\varphi_{\theta})$ the (see (8) and (11), respectively) reduce to
$\pi^{\varphi_{\theta}}(x,y)=\rho_{0}(x)\,e^{y^{\top}m_{1}(x)/(\varepsilon\sigma_{1}^{2})}v_{\theta}(y)\big/\penalty 50c_{\theta}(x)$ (20)
$c_{\theta}(x)=\int\limits_{\mathbb{R}^{d}}e^{y^{\top}m_{1}(x)/(\varepsilon\sigma_{1}^{2})}v_{\theta}(y)\,{\mathrm{d}}y=\sum\limits_{k=1}^{K}\widetilde{\alpha}_{k}(x)$ (21)
and
$\widetilde{\alpha}_{k}(x)=\alpha_{k}\exp\left\{\frac{r_{k}^{\top}m_{1}(x)}{\varepsilon\sigma_{1}^{2}}+\frac{\|S_{k}^{1/2}m_{1}(x)\|^{2}}{2\varepsilon\sigma_{1}^{2}}\right\}.$ (22)
The log-potential estimate $\widehat{\varphi}$ is constructed through minimization of the empirical risk
over all admissible values of $\theta$. We call the described algorithm LightSB-OU and provide its pseudocode in Appendix F.2. The standard LightSB algorithm (Korotin et al., 2024) performs the same steps but replaces $\mathsf{q}_{1}(y\,|\,x)$ with the Gaussian kernel $(2\pi\varepsilon)^{-d/2}e^{-\|y-x\|^{2}/(2\varepsilon)}$ corresponding to the transition density of the scaled Wiener process (3) with $\sigma=\sqrt{\varepsilon}$.
Below we thoroughly evaluate our solver on setups with both synthetic (Section 4.1) and real data sets (Sections 4.2 and 4.3). The method demonstrates a stable increase in quality compared to the original LightSB and shows itself well in comparison with other Schrödinger bridge solvers. The source code is available at GitHub [^2] https://github.com/denvar15/Tight-Bounds-for-Schrodinger-Potential-Estimation-in-Unpaired-Data-Translation.
4.1 Evaluation on the Gaussian Mixture
We start with evaluation of the modified LightSB algorithm on a Gaussian mixture problem. For this experiment, we used mixtures of 25 Gaussians under three configurations: uniform grid with standard covariance matrices, random grid locations with standard covariance matrices, uniform grid with anisotropic random covariance matrices, with the sliced approximation of the Wasserstein distance $\mathbb{W}_{1}$ as the metric, see Appendix F.2 for definition of metrics.
Table 3 summarizes the performance for $K=30$ potential approximation components. With a sufficient number of potential approximation components, our approach demonstrates consistent improvements in both $\mathbb{W}_{1},$ MMD and Mode Coverage (see Appendix F.3 for detailed plots and Appendix F.6 for hyperparameters values used).
Table 2: Comparison of the LightSB and LightSB-OU performance for $K=30$ across experiments with Gaussian mixtures with $25$ components
Experiment Metric LightSB LightSB-OU
Standard Sliced $\mathbb{W}_{1}$ $0.260\pm 0.016$ $\mathbf{0.156\pm 0.018}$
[…]
Mode Coverage $24.2\pm 0.4$ $\mathbf{25.0\pm 0.0}$
Irregular Sliced $\mathbb{W}_{1}$ $0.525\pm 0.024$ $\mathbf{0.214\pm 0.028}$
[…]
Mode Coverage $21.8\pm 0.4$ $\mathbf{25.0\pm 0.0}$
Anisotropic Sliced $\mathbb{W}_{1}$ $0.255\pm 0.017$ $\mathbf{0.206\pm 0.024}$
[…]
Mode Coverage $24.4\pm 0.5$ $\mathbf{25.0\pm 0.0}$
Table 3: The quality of intermediate distribution restoration for the single-cell data
Solver $\mathbb{W}_{1}$ metric
OT-CFM (Tong et al., 2024a) $0.790\pm 0.068$
$\mathrm{SF}]^{2}$M-Exact (Tong et al., 2024b) $0.793\pm 0.066$
LightSB-OU (ours) $\mathbf{0.812\pm 0.020}$
LightSB (Korotin et al., 2024) $0.823\pm 0.017$
Reg. CNF (Finlay et al., 2020) $0.825\pm\text{N/A}$ [^3] The authors did not report the standard deviation.
T. Net (Tong et al., 2020) $0.848\pm\text{N/A}$3
DSB (De Bortoli et al., 2021) $0.862\pm 0.023$
I-CFM (Tong et al., 2024a) $0.872\pm 0.087$
$\mathrm{SF}]^{2}$M-Geo (Tong et al., 2024b) $0.879\pm 0.148$
NLSB (Koshizuka & Sato, 2023) $0.970\pm\text{N/A}$3
$\mathrm{SF}]^{2}$M-Sink (Tong et al., 2024b) $1.198\pm 0.342$
SB-CFM (Tong et al., 2024a) $1.221\pm 0.380$
DSBM (Shi et al., 2023) $1.775\pm 0.429$
4.2 Evaluation on the Single Cell Data
To comprehensively evaluate LightSB‑OU, we conducted experiments on biological data. Following Tong et al. (2020), we considered the problem of transporting a cell distribution from time $t_{i-1}$ to time $t_{i+1}$ for $i\in\{1,2,3\}$. The results of Tong et al. (2024b) allow us to compare the performance of LightSB-OU not only with LightSB but also with many other methods (see Table 3). We predicted the cell distribution at the intermediate time $t_{i}$ and computed the Wasserstein distance $\mathbb{W}_{1}$ between the predicted distribution and the ground truth. The results were averaged across three setups ($i=1,2,3$). To ensure statistical robustness, we repeated the experiment five times. Using the default LightSB parameters from the original paper (Korotin et al., 2024), the baseline achieved $\mathbb{W}_{1}=0.823\pm 0.017$ and was outperformed by our OU-modified version ($\mathbb{W}_{1}=0.810\pm 0.020$). The hyperparameter values of LightSB-OU are reported in Appendix F.6.
4.3 Unpaired Image-to-Image Translation
(a) LightSB: FID = 24.1
(b) LightSB-OU: FID = 24.0
Figure 2: Translation results from the latent space of ALAE using LightSB (left) and LightSB-OU(right). Top: Adult to Child – skin color and face shape preserved. Middle: Male to Female – glasses and face shape are preserved. Bottom: Male to Female – skin color and face shape preserved.
To further illustrate the importance of the reference process for data generation, we consider a task of unpaired image-to-image translation. Following the experimental setup of Korotin et al. (2024), we employ the ALAE autoencoder (Pidhorskyi et al., 2020) and focus on the Adult-to-Child translation task using the full $1024\times 1024$ FFHQ dataset (Karras et al., 2019).
Below, we present comparative translation results for selected samples in Figure 2. While the baseline LightSB successfully performs the translation, it sometimes fails to preserve key attributes of the source distribution such as skin tone, eye color, and facial structure leading to noticeable deviations. In contrast, LightSB-OU achieves more accurate translations while better retaining the distinctive features of the original distribution. We emphasize that these examples are not intended to suggest universal superiority but rather to showcase specific strengths of our method in certain cases (the hyperparameter values are specified in Appendix F.6).
5 Conclusion
In the present paper, we established that the empirical risk minimizer $\widehat{\pi}$ defined in (12) satisfies a Bernstein-type bound (see Theorem 1). This is a consequence of exponential ergodicity of the OU process and of some properties of the empirical risk (13) exhibiting a kind of curvature when $bT\gtrsim\log\log n$. The importance of the reference process was also illustrated with numerical experiments on artificial and real-world data. Further research may include an extension of Theorem 1 to the case when $\rho_{0}$ has an unbounded support and to heavy-tailed distributions $\rho_{T}$. It is also of particular interest whether the result of Theorem 1 holds if one considers a larger class of exponentially ergodic reference processes.
Reproducibility statement
The provide the proof of our main result, Theorem 1, in Appendix B. In Appendix B.1, we discuss some insights behind the proof of Theorem 1 and then perform rigorous derivations in Appendix B.2. The proofs of all the auxiliary statements are collected in Appendix.
In Appendix F, we provide additional details for the numerical experiments described in Section 4 including the pseudocode of the LightSB-OU algorithm (see Algorithm 1). The code for the numerical experiments is available in the supplementary material and at GitHub [^4] https://github.com/denvar15/Tight-Bounds-for-Schrodinger-Potential-Estimation-in-Unpaired-Data-Translation.
We have not used large language models (LLMs) for retrieval, discovery or research ideation. All the mathematical derivations and numerical experiments were performed solely by the authors. We used LLMs just to polish writing.
Acknowledgements
The work was supported by the grant for research centers in the field of AI provided by the Ministry of Economic Development of the Russian Federation in accordance with the agreement  [^00] C313925P4E0002 and the agreement with HSE University №139-15-2025-009.
Appendix A On Assumption 4
@@
In Section 3, we suppose that the log-potential $\varphi^{*}$ associated with the optimal coupling $\pi^{*}$ meets Assumption 4. While the requirement $\varphi^{*}(x)\leqslant M$ is mild, the lower bound $\varphi^{*}(x)\geqslant-L\|\Sigma^{-1/2}(x-m)\|^{2}-M$ needs a discussion. We are going to provide a sufficient condition for Assumption 4 to hold. Let us define $\varphi_{\max}^{*}=\sup\limits_{x\in\mathbb{R}^{d}}\varphi^{*}(x)<+\infty$. In view of (9), we have
@@
recently considered by Yakovlev & Puchkin (2025) in the context of generative diffusion models. Here $g:[0,1]^{p}\rightarrow\mathbb{R}^{d}$ is a smooth map defined on a low-dimensional latent space. In (Yakovlev & Puchkin, 2025), the authors discussed that such a model is quite general and suitable for description of high-dimensional data occupying a vicinity of a smooth low-dimensional manifold.
Appendix B Proof of Theorem 1
B.1 Proof Insights
Finally, we would like to discuss some intuition behind the proof of Theorem 1. Rigorous derivations are deferred to Appendix B. As we discussed in Section 3, Assumptions 2, 3, and 4 allow us to use concentration inequalities for sub-exponential random variables and quantify the deviation of the empirical risks $\widehat{\mathcal{L}}(\varphi)$, $\varphi\in\Phi$, from its expectations. However, to get rates of convergence possibly sharper than ${\mathcal{O}}(1/\sqrt{n})$, we must verify a Bernstein-type condition. That is, we have to show that both